If you have a garden that is 5mx5mx5mx5m and is 50cm deep. how do you work out the volume of soil you need?
Answer by mananth(16946)

Volume of mud = L * W * H
=5*5*0.5 ( 50cm = 0.5m)
=12.5 cu.m